OpenStudy (anonymous):

If angles 1 and 2 are a linear pair, and m<1 = 3x + 12 and m<2 = x + 44, find the measures of these two angles

OpenStudy (blurbendy):

3x + 12 = x + 44 2x = 32 x = 16 degrees.

OpenStudy (blurbendy):

Then put x back in to both equations and find the measures of both the angles.

OpenStudy (sirm3d):

linear pair means the sum of the angles is 180. (3x + 12) + (x + 44) = 180 solve x.

OpenStudy (sirm3d):

don't forget to sub the value in (3x + 12) and (x + 44) to get the measures of the two angles.

OpenStudy (blurbendy):

ah, ty for the correction

OpenStudy (anonymous):

x=31